# F-LMM: Grounding Frozen Large Multimodal Models

### v1

@@ Abstract
Endowing Large Multimodal Models (LMMs) with visual grounding capability can significantly enhance AIs’ understanding of the visual world and their interaction with humans. However, existing methods typically fine-tune the parameters of LMMs to learn additional segmentation tokens and overfit grounding and segmentation datasets. Such a design would inevitably cause a catastrophic diminution in the indispensable conversational capability of general AI assistants. In this paper, we comprehensively evaluate state-of-the-art grounding LMMs across a suite of multimodal question-answering benchmarks, observing pronounced performance drops that indicate vanishing general knowledge comprehension and weakened instruction following ability. To address this issue, we present F-LMM—grounding frozen off-the-shelf LMMs in human-AI conversations—a straightforward yet effective design based on the fact that word-pixel correspondences conducive to visual grounding inherently exist in the attention weights of well-trained LMMs. Using only a few trainable CNN layers, we can translate word-pixel attention weights to mask logits, which a SAM-based mask refiner can further optimise. Our F-LMM neither learns special segmentation tokens nor utilises high-quality grounded instruction-tuning data, but achieves competitive performance on referring expression segmentation and panoptic narrative grounding benchmarks while completely preserving LMMs’ original conversational ability. Additionally, with instruction-following ability preserved and grounding ability obtained, our F-LMM can perform visual chain-of-thought reasoning and better resist object hallucinations. Code and models will be released at https://github.com/wusize/F-LMM.
Figure 1: An example of user-AI conversation around an image. Left: The current state-of-the-art grounding model GLaMM [57] is effective for grounded conversation when prompted by "answer with interleaved masks", but fails to follow user instruction to answer a single word (yes or no) and misunderstands the question as a referring segmentation prompt. Right: Our F-LMM preserves instruction-following ability while being able to perform visual grounding.
@@ 1 Introduction
As one of the pivotal milestones in AGI, recent Large Multimodal Models (LMMs)—integrating Large Language Models (LLMs) with visual signals—have demonstrated remarkable success in multimodal understanding, reasoning and interaction [39, 37, 38, 42, 32, 62, 75]. To further advance LMMs with better perception capability, a recent line of research [77, 28, 57, 58, 67, 80] that visually grounds language contents in user-model conversations has drawn increasing attention. This explicit association between key phrases or words with visual objects greatly enhances LMMs’ understanding of the visual world and allows for more intuitive and meaningful human-AI interactions.
@@
By design, one commonly adopted build (Figure 2(b)) for visually grounding language contents is connecting LMMs with a mask head (e.g., Segment Anything Model (SAM) [26]), wherein both the LLM backbone and the mask head are fine-tuned with well-prepared visual grounding data that contains segmentation annotations. Also, some additional learnable tokens (e.g., [SEG]) are introduced to the LMMs’ vocabulary, to directly associate key phrases or words with visual objects in conversations. However, this design will inevitably provoke a catastrophic diminution in general knowledge comprehension and instruction-following ability due to the following reasons. First, existing segmentation and visual grounding data only contain elementary patterns for answering simple grounding prompts. Second, during the fine-tuning stage, the LMMs are mainly optimised for effectively modelling the relationship between key phrases or words and special segmentation tokens, i.e., overfitting the segmentation and grounding data. Therefore, the conversational ability that is indispensable in building general AI assistants is sacrificed. For instance, the state-of-the-art grounding model GLaMM [57] fails to answer a simple yes-or-no question (Figure 1). Moreover, quantitative evaluations of existing grounding LMMs in conversational ability are presented in Table 1, with zero or near-zero scores on general multimodal question-answering benchmarks necessitating instruction-following ability.
@@
To deal with this dilemma, one possible option is to collect high-quality training data encompassing both meaningful conversations and mask annotations. For example, LLaVA-G [77] annotates the 150k LLaVA-Instruct data samples [39] with segmentation masks so that the LMMs simultaneously learn to chat and segment. Nonetheless, annotating high-quality grounded conversation data is costly and hard to scale up. Despite being trained on the costly annotated data, LLaVA-G still lags behind general-purpose LMMs on multimodal understanding tasks. Furthermore, training on large-scale annotated data normally consumes significant computational resources, which is, obviously not a resource-efficient solution.
@@
In this paper, we propose a straightforward yet effective design, i.e., grounding frozen LMMs (dubbed as F-LMM) in human-AI conversations. Thinking from first principles, we argue that freezing the parameters of well-trained LMMs is the most feasible design choice for completely maintaining the original excellent conversational ability when building general-purpose grounding LMMs. In particular, we take inspiration from the built-in interpretability of the attention mechanism in transformers [65, 6] that represents interrelations between text and image contents in design. We observe that off-the-shelf LMMs already produce word-pixel correspondences necessary for visual grounding, despite not being directly pre-trained with region or pixel annotations. As exhibited in Figure 2(a), we visualise word-image attention maps from frozen LMMs via K-Means clustering, demonstrating prominent geometric and spatial cues of the objects.  [^1] For better visibility, we perform K-Means clustering on the stack-up of all attention maps collected in a forward pass instead of selecting a single attention map. For example, coarse visual grounding masks for key phrases (e.g., "two girls", "two elephants", and "the plate") in language sentences can emerge from attention maps in LMMs. Therefore, our F-LMM takes these visual-language correspondences as useful segmentation priors for decoding grounding masks, neither further tuning the LMMs’ weights nor learning a special segmentation token to model object locations, as shown in Figure 2(c).
@@
Notably, the only trainable part of our F-LMM is a mask head. It comprises a CNN-based mask decoder (a tiny U-Net [59]) that translates stacked attention maps to mask logits and a light-weight mask refiner (retrofitted from SAM [26]’s mask head) that uses additional image and language cues to refine the semantic-agnostic masks from the mask decoder. Moreover, we only use the RefCOCO(+/g) [23, 44] and PNG [17] datasets as our training data, enabling LMMs to segment user-described objects and ground key phrases or words in a text sequence. Unlike previous works [77, 57, 58], our F-LMM eliminates the necessity for high-quality conversation data that are annotated with masks to preserve conversational ability when learning grounding.
@@
In experiments, our F-LMM maintains the original excellence of off-the-shelf LMMs on general question-answering benchmarks, while achieving competitive results on referring segmentation and phrase grounding. Compared with existing grounding LMMs, F-LMM offers the best balance between grounding and chat capabilities. In addition, with instruction-following ability preserved and grounding ability obtained, F-LMM unleashes visual chain-of-thought reasoning in a zero-shot manner by exhibiting improvement on the VisCoT benchmark [60] and better resists object hallucinations on the POPE benchmark [34].
@@ 2 Related Work
Large Multimodal Models. Recent advancements in LMMs  [2, 31, 14, 39, 37, 38, 72, 3, 30, 35, 45, 42, 32, 62, 29] have been fueled by the success of LLMs [4, 1, 78, 63, 64, 22, 12, 49, 46] since the debut of GPT series [54, 55, 4, 1] that feature an auto-regressive framework based on transformer decoder [65]. These LLMs possess general world knowledge and excellent conversational ability to follow human instructions, thanks to large-scale generative pre-training [4] and supervised finetuning on instruction-tuning data [68] or human feedback [48]. By integrating image representations from vision encoders [53, 76] to LLMs, LMMs enable visual understanding and reasoning as AI assistants. This integration is usually established by a multilayer perceptron (MLP) that directly maps image features to the LLMs’ input embedding space [39, 37, 38, 42, 32, 62, 75] or a cross-attention module that abstracts the image contents with a set of query embeddings [2, 31, 3, 72]. In our research, we build F-LMM on LMMs of the former type (MLP-based), which preserves images’ 2-D topological structure in the cross-modal integration.
@@
Visual Segmentation. The task of predicting 2D masks for visual objects is known as image segmentation, which can be categorised into semantic segmentation [8, 5, 81, 11], instance segmentation [20, 10, 79] and panoptic segmentation [25, 9, 70, 33] depending on whether the goal is to differentiate pixel semantics or object instances. These standard segmentation approaches rely on a pre-defined set of object classes for recognition. In contrast, referring expression segmentation (RES) [23, 44, 47, 82, 43, 71, 36] involves segmenting objects based on free-form human language descriptions, allowing for enhanced human-model interaction. Additionally, panoptic narrative grounding (PNG) [17, 15, 66, 18] requires segmenting masks for key phrases or words in a sentence. In this study, we leverage RES and PNG tasks to evaluate the grounding capability of LMMs. In addition, the prompt-based SAM [26] pre-trained on billion-scale high-quality mask data have become a constituent component in many grounding LMMs to boost segmentation performance. We also adopt SAM’s mask head to initialise our mask refiner.
@@
Grounding Large Multimodal Models. Grounding Large Multimodal Models [50, 7, 3, 73, 77, 80, 58, 57, 67, 28, 69, 51, 75] can localise language contents during user-model conversations. Some approaches [50, 7, 73, 3] represent coordinates of bounding boxes as texts and train LMMs to predict the coordinates in a generative manner. Several recent works [28, 77, 58, 67, 57, 51] train LMMs to predict a special segmentation token for encoding the grounded object and utilise a segmentation head (e.g., SAM [26]) to decode object masks. This study mainly focuses on grounding LLMs with segmentation ability for visual perception. To obtain competitive visual grounding performance, existing works extensively fine-tune the parameters of LMMs on a large amount of segmentation [81, 5, 25, 56, 19] and grounding [23, 44, 24, 27, 52, 17] datasets. And to balance the LMMs’ grounding and conversational abilities, there are efforts [77, 58, 80] to collect high-quality instruction-tuning data annotated with segmentation masks. In contrast, we make the first attempt to build grounding LMMs on top of off-the-shelf LMMs without fine-tuning their parameters. Furthermore, we bypass the need for grounded instruction-tuning data to preserve decent chat ability.
@@ 3 Method
Figure 3: The overall pipeline of F-LMM. The word-image attention maps from the frozen LLM serve as segmentation priors for the mask head. The mask head encompasses a mask decoder that translates attention weights to mask logits and a mask refiner that optimises the mask decoder’s predictions. $M$ and $N$ represent the numbers of transformer layers and attention heads.
In this section, we introduce our F-LMM by first probing the causal attention mechanism in LMMs with visualisations of word-image attention maps in Sec 3.1. Then, we elaborate on F-LMM exploiting segmentation priors from frozen LMMs for visual grounding using the mask head in Sec 3.2. Finally, we show how to perform referring expression segmentation and phrase grounding with our F-LMM for human-AI conversations in Sec 3.3. The overall pipeline is illustrated in Figure 3.
@@ 3.1 Segmentation Priors from Frozen LMM
Vision-Language Sequence. A typical build of a Large Multimodal Model (LMM)  [^2] In this paper, the term ‘multimodal’ stands for vision and language modalities. comprises an image encoder $f_{v}$ (e.g., CLIP [53]  [^3] The image encoder might be any vision model that is pre-trained on image-text pairs. We use the classic term ‘CLIP’ in this paper to represent all such models for brevity.), a vision-language projector $f_{p}$, and a Large Language Model (LLM) $f_{\mathrm{llm}}$. The inputs to an LMM are usually an image ${\mathbf{X}}_{v}\in{\mathbb{R}}^{3\times H\times W}$ and the associated text ${\mathbf{X}}_{t}$. The input image is first encoded by the vision encoder $f_{v}$ and then mapped to the input space of the LLM $f_{\mathrm{llm}}$ by the projector $f_{p}$:
${\mathbf{Z}}_{v}=f_{p}(\texttt{Flatten}(f_{v}({\mathbf{X}}_{v})))\in{\mathbb{R}}^{hw\times d},$ (1)
@@
where $h$ and $w$ are the height and width of projected feature maps via $f_{v}$. The Flatten operation unfolds the 2-D image feature map to a 1-D sequence. The constant $d$ is the hidden state dimension of the LLM $f_{\mathrm{llm}}$. Likewise, the text input is first encoded as discrete tokens and then mapped to text embeddings:
${\mathbf{Z}}_{t}=\texttt{Embed}(\texttt{Tokenize}({\mathbf{X}}_{t}))\in{\mathbb{R}}^{L\times d},$ (2)
@@
where $L$ denotes the length of text embeddings. The visual-language sequence input to the LLM $f_{\mathrm{llm}}$ is a concatenation of image and text embeddings: ${\mathbf{Z}}=\{{\mathbf{Z}}_{v},{\mathbf{Z}}_{t}\}\in{\mathbb{R}}^{(hw+L)\times d}$.
Segmentation Priors in Self-Attention. The vision-language sequence is mainly processed by causal self-attentions [65, 54] in the LLM $f_{\mathrm{llm}}$, including inner product and weighted-sum operations. Specifically, for a word token with position index $i$ in the vision-language sequence ${\mathbf{Z}}$ , its embedding ${\bm{z}}^{i}$ is updated by the weighted sum of the first $i$ embeddings: $\hat{{\bm{z}}}^{i}=\texttt{SoftMax}(\frac{{\bm{z}}^{i}\cdot{\mathbf{Z}}[:i]}{d})\cdot{\mathbf{Z}}[:i]$, where $\texttt{SoftMax}(\frac{{\bm{z}}^{i}\cdot{\mathbf{Z}}[:i]}{d})$ is the attention weights. Here, we omit the residual layers and feedforward layers for brevity.
Figure 4: Visualisations of word-image attention maps. The letters $m$ and $n$ indicate that the attention map is derived from the $n$-th attention head of the $m$-th transformer layer. Though noisy, the objects are observable in the attention maps. The visibility is further enhanced when we stack up all the attention maps and perform KMeans clustering.
$\bm{a}^{i}=\texttt{Unflatten}(\texttt{SoftMax}(\frac{{\bm{z}}^{i}\cdot{\mathbf{Z}}[:i]}{d})[:hw])\in{\mathbb{R}}^{h\times w},$ (3)
where Unflatten restores the 2-D spatial structure from the 1-D sequence to form an attention map. In Figure 4, we visualise such word-image attention maps from various transformers layers and attention heads in an LMM (i.e., DeepseekVL-1.3B [42]). The objects’ shape and location can be observed in word-image attention maps of certain layers or heads. The visibility is further enhanced when we stack the attention maps from all layers and heads and perform K-Means clustering. It can be observed that the attention maps offer meaningful segmentation priors with spatial and geometric cues for grounding objects visually.
3.2 Visual Grounding with Mask Head
Mask Decoder. The mask decoder $f_{d}$ is a 2-D CNN model that transforms the word-image attention maps of grounded objects into mask logits, which is instantiated by a 3-stage U-Net [59]. Please refer to Sec A.2 of the appendix for details of the mask decoder. The extraction of word-image attention map $\bm{a}^{i}$ for a word token with position index $i$ is illustrated in Eq. 3 and Figure 3. For an object described by multiple words, we merge their corresponding word-image attention maps to a single attention map $\bm{a}$ via element-wise average or max operation. The attention map $\bm{a}$ is further normalised as $\bm{a}/\mathrm{sum}(\bm{a})$ so that all elements sum to 1. Considering $M$ layers and $N$ attention heads, we stack the $MN$ attention maps as $\bm{A}\in{\mathbb{R}}^{MN\times h\times w}$, which forms the input to a mask decoder. Given the importance of high input resolution for segmentation models, we upsample the stacked attention maps $\bm{A}$ to $h^{\prime}\times w^{\prime}$ by bilinear interpolation before feeding it to a mask decoder, where $h^{\prime}>h$ and $w^{\prime}>w$. In practice, we set $h^{\prime}=w^{\prime}=64$. Then, the mask decoder maps $\bm{A}$ into mask logits: ${\mathbf{M}}_{\mathrm{logits}}=f_{d}(\bm{A})$. We derive the corresponding binary mask via ${\mathbf{M}}_{\mathrm{pred}}={\mathbf{M}}_{\mathrm{logits}}>0$. During training, the mask decoder is optimised with BCE and DICE losses [61].
3.3 User-AI Interaction with Grounding
We elaborate on how F-LMM works for user-AI conversations in two typical scenarios, i.e., referring expression segmentation and phrase grounding.
Referring Expression Segmentation. In this scenario, the model is supposed to segment user-described objects. Existing works [28, 57] request the LMM generate a special segmentation token in the answer, which is then decoded as the mask of the described object. In our F-LMM, we can directly perform grounding for user descriptions using word-image attention maps and text embeddings.
Phrase Grounding. In user-model conversations, the grounding LMMs can localise key phrases or words when chatting with the user. Unlike existing works [58, 57] that generate special tokens to indicate grounded objects, we use the spaCy toolkit [21] to obtain object words and phrases in the texts and decode grounding mask from frozen LMMs. This disentanglement of text generation and phrase selection in our design also allows users to decide which words or phrases to ground.
4 Experiments
@@
Table 1: The main evaluation results on question-answering benchmarks, referring expression segmentation (RES) benchmark and panoptic narrative grounding (PNG) benchmark. MMB: MMBench; LLaVAW: LLaVA-In-the-Wild; RefC(+/g): RefCOCO(+/g). LLaVA-1.6 and MGM-HD take high-resolution image inputs. LLaVA-1.6-M-7B means the model is based on Mistral-7B [22]. GLaMM-FS-7B means we use the ‘FullScope’ version of GLaMM.
@@
MME MMB MMVet LLaVAW RefC RefC+ RefCg All Thing Stuff
Specialised Segmentation Models
MCN [43] - - - - 62.4 50.6 49.2 54.2 48.6 61.4
LAVT [71] - - - - 72.7 62.1 61.2 - - -
GRES [36] - - - - 73.8 66.0 65.0 - - -
X-Decoder [82] - - - - - - 64.6 - - -
SEEM [83] - - - - - - 65.7 - - -
PNG [17] - - - - - - - 55.4 56.2 54.3
PPMN [15] - - - - - - - 59.4 57.2 62.5
XPNG [18] - - - - - - - 63.3 61.1 66.2
Existing Grounding LMMs

### v2

@@ Abstract
Endowing Large Multimodal Models (LMMs) with visual grounding capability can significantly enhance AIs’ understanding of the visual world and their interaction with humans. However, existing methods typically fine-tune the parameters of LMMs to learn additional segmentation tokens and overfit grounding and segmentation datasets. Such a design would inevitably cause a catastrophic diminution in the indispensable conversational capability of general AI assistants. In this paper, we comprehensively evaluate state-of-the-art grounding LMMs across a suite of multimodal question-answering benchmarks, observing drastic performance drops that indicate vanishing general knowledge comprehension and weakened instruction following ability. To address this issue, we present F-LMM—grounding frozen off-the-shelf LMMs in human-AI conversations—a straightforward yet effective design based on the fact that word-pixel correspondences conducive to visual grounding inherently exist in the attention mechanism of well-trained LMMs. Using only a few trainable CNN layers, we can translate word-pixel attention weights to mask logits, which a SAM-based mask refiner can further optimise. Our F-LMM neither learns special segmentation tokens nor utilises high-quality grounded instruction-tuning data, but achieves competitive performance on referring expression segmentation and panoptic narrative grounding benchmarks while completely preserving LMMs’ original conversational ability. Additionally, with instruction-following ability preserved and grounding ability obtained, F-LMM can be directly applied to complex tasks like reasoning segmentation, grounded conversation generation and visual chain-of-thought reasoning. Our code can be found at https://github.com/wusize/F-LMM.
Figure 1: An example of user-AI conversation around an image. Left: The current state-of-the-art grounding model GLaMM [60] is effective for grounded conversation when prompted by ”answer with interleaved masks”, but fails to follow user instruction to answer a single word (yes or no) and misunderstands the question as a referring segmentation prompt. Right: Our F-LMM preserves instruction-following ability while being able to perform visual grounding.
@@ 1 Introduction
Large Multimodal Models (LMMs), which integrate Large Language Models (LLMs) with visual signals, have demonstrated remarkable success in multimodal understanding, reasoning and interaction [39, 38, 40, 44, 34, 65, 78]. To further advance LMMs with better perception capability, a recent line of research [80, 28, 60, 61, 70, 84] that visually grounds language contents in user-model conversations has drawn increasing attention. This explicit association between key phrases/words and visual objects greatly enhances LMMs’ understanding of the visual world and allows for more intuitive and meaningful human-AI interactions.
@@
By design, one commonly adopted build (Figure 2(b)) for visually grounding language contents is connecting LMMs with a mask head (e.g., Segment Anything Model (SAM) [26]), wherein both the LLM backbone and the mask head are fine-tuned with well-prepared visual grounding data that contains segmentation annotations. Also, some additional learnable tokens (e.g., [SEG]) are introduced to the LMMs’ vocabulary, to directly associate key phrases or words with visual objects in conversations. However, this design will inevitably provoke a catastrophic diminution in general knowledge comprehension and instruction-following ability due to the following reasons. First, existing segmentation and visual grounding data only contain elementary patterns for answering simple grounding prompts. Second, during the fine-tuning stage, the LMMs are mainly optimised for effectively modelling the relationship between key phrases or words and special segmentation tokens, i.e., overfitting the segmentation and grounding data. Therefore, the conversational ability is sacrificed. For instance, the state-of-the-art grounding model GLaMM [60] fails to answer a simple yes-or-no question (Figure 1). Moreover, quantitative evaluations of existing grounding LMMs in conversational ability are presented in Table 1, with zero or near-zero scores on general multimodal question-answering benchmarks necessitating instruction-following ability.
@@
One possible option to deal with this dilemma is to collect high-quality training data encompassing both meaningful conversations and mask annotations. For example, LLaVA-G [80] annotates the 150k LLaVA-Instruct data samples [39] with segmentation masks so that the LMMs simultaneously learn to chat and segment. Nonetheless, annotating high-quality grounded conversation data is costly and hard to scale. Despite being trained on costly annotated data, LLaVA-G still lags behind general-purpose LMMs on multimodal understanding tasks. Furthermore, training on large-scale annotated data normally consumes significant computational resources, which is, obviously not a resource-efficient solution.
@@
In this paper, we propose a simple yet effective design, i.e., grounding frozen LMMs (dubbed as F-LMM) in human-AI conversations. We argue that freezing the parameters of well-trained LMMs is the most practical design choice for fully preserving the original excellent conversational ability when building general-purpose grounding LMMs. In particular, we take inspiration from the built-in interpretability of the attention mechanism in transformers [68, 6] that represents interrelations between text and image contents in design. We observe that off-the-shelf LMMs already produce word-pixel correspondences necessary for visual grounding, despite they were not explicitly pre-trained with region or pixel annotations. As illustrated in Figure 2(a), we visualise word-image attention maps from frozen LMMs via K-Means clustering, revealing notable geometric and spatial cues of the objects  [^1] For better visibility, we perform K-Means clustering on the stack-up of all attention maps collected in a forward pass instead of selecting a single attention map.. For example, coarse visual grounding masks for key phrases (e.g., “two girls”, “two elephants”, and “the plate”) in language sentences emerge from attention maps in LMMs. Therefore, our F-LMM takes these visual-language correspondences as useful segmentation priors for decoding grounding masks, without further tuning the LMMs’ weights or learning a special segmentation token to model object locations, as shown in Figure 2(c).
@@
The only trainable part of our F-LMM is a mask head plus a keyword selector. The mask head comprises a CNN-based mask decoder (a tiny U-Net [62]) that translates stacked attention maps to mask logits and a light-weight mask refiner (retrofitted from SAM [26]’s mask head) that uses additional image and language cues to refine the semantic-agnostic masks from the mask decoder. The keyword selector is a linear layer that discovers object nouns in text sequences, automating the process of grounded human-AI conversation. Moreover, we only use the RefCOCO(+/g) [24, 46] and PNG [16] datasets as our training data, enabling LMMs to segment user-described objects and ground key phrases or words in a text sequence. Unlike previous works [80, 60, 61], our F-LMM eliminates the necessity for high-quality conversation data that are annotated with masks to preserve conversational ability when learning grounding.
@@
Our experiments demonstrate that F-LMM maintains the original excellence of off-the-shelf LMMs on general question-answering benchmarks, while achieving competitive results on referring segmentation and phrase grounding. In more complex tasks like reasoning segmentation, grounded conversation generation and visual chain-of-thought reasoning, F-LMM achieves better or comparable results when contrasted with models specially trained for such tasks. Compared with existing grounding LMMs, F-LMM offers the best balance between grounding and chat capabilities.
@@ 2 Related Work
Large Multimodal Models. Recent advancements in LMMs  [2, 31, 13, 39, 38, 40, 75, 3, 30, 36, 47, 44, 34, 65, 29] have been fueled by the success of LLMs [4, 1, 81, 66, 67, 22, 12, 51, 48] since the debut of GPT series [56, 57, 4, 1] that feature an auto-regressive framework based on transformer decoders [68]. These LLMs possess general world knowledge and excellent conversational ability to follow human instructions, thanks to large-scale generative pre-training [4] and supervised finetuning on instruction-tuning data [71] or human feedback [50]. By integrating image representations from vision encoders [58, 79] to LLMs, LMMs enable visual understanding and reasoning in AI assistants. This integration is usually established by a multilayer perceptron (MLP) that directly maps image features to the LLMs’ input embedding space [39, 38, 40, 44, 34, 65, 78] or a cross-attention module that abstracts the image contents with a set of query embeddings [2, 31, 3, 75]. In our research, we build F-LMM on LMMs of the former type (MLP-based), which preserves images’ 2-D topological structure in the cross-modal integration.
@@
Visual Segmentation. The task of predicting 2D masks for visual objects is known as image segmentation, which can be categorised into semantic segmentation [8, 5, 87, 10], instance segmentation [19, 11, 83] and panoptic segmentation [25, 9, 73, 32] depending on whether the goal is to differentiate pixel semantics or object instances. These standard segmentation approaches rely on a pre-defined set of object classes for recognition. In contrast, referring expression segmentation (RES) [24, 46, 49, 88, 45, 74, 37] involves segmenting objects based on free-form human language descriptions, allowing for enhanced human-model interaction. Additionally, panoptic narrative grounding (PNG) [16, 14, 69, 17] requires segmenting masks for key phrases or words in a sentence. In this study, we mainly leverage RES and PNG tasks to evaluate the grounding capability of LMMs. Besides, we also test LMMs’ segmentation ability in complex scenarios that necessitate reasoning [60, 28, 63]. Moreover, the prompt-based SAM [26] pre-trained on billion-scale high-quality mask data has become a constituent component in many grounding LMMs to boost segmentation performance. We also adopt SAM’s mask head to initialise our mask refiner.
@@
Grounding Large Multimodal Models. Grounding Large Multimodal Models [53, 7, 3, 76, 80, 84, 61, 60, 70, 28, 72, 54, 78, 85] can localise language contents during user-model conversations. Some approaches [53, 7, 76, 3] represent coordinates of bounding boxes as texts and train LMMs to predict the coordinates in a generative manner. Several recent works [28, 80, 61, 70, 60, 54] train LMMs to predict a special segmentation token for encoding the grounded object and utilise a segmentation head (e.g., SAM [26]) to decode object masks. This study mainly focuses on grounding LMMs with segmentation ability for visual perception. To obtain competitive visual grounding performance, existing works extensively fine-tune the parameters of LMMs on a large amount of segmentation [87, 5, 25, 59, 18] and grounding [24, 46, 23, 27, 55, 16] datasets. And to balance the LMMs’ grounding and conversational abilities, there are efforts [80, 61, 84] to collect high-quality instruction-tuning data annotated with segmentation masks. In contrast, we make the first attempt to build grounding LMMs on top of off-the-shelf LMMs without fine-tuning their parameters. Furthermore, we bypass the need for grounded instruction-tuning data to preserve decent chat ability.
@@ 3 Method
Figure 3: The overall pipeline of F-LMM. The word-image attention maps from the frozen LLM serve as segmentation priors for the mask head. The keyword selector discovers object nouns in the text sequence. The mask head encompasses a mask decoder that translates attention weights to mask logits and a mask refiner that optimises the mask decoder’s predictions. $M$ and $N$ represent the numbers of transformer layers and attention heads.
Figure 4: Visualisations of word-image attention maps. The letters $m$ and $n$ indicate that the attention map is derived from the $n$-th attention head of the $m$-th transformer layer.
In this section, we introduce our F-LMM by first probing the causal attention mechanism in LMMs with visualisations of word-image attention maps in Sec 3.1. Then, we elaborate on F-LMM exploiting segmentation priors from frozen LMMs for visual grounding using the mask head in Sec 3.2. Finally, we show how to automate the process of grounded conversation with a linear keyword selector to indicate words of grounding targets in Sec 3.3. The overall pipeline is illustrated in Figure 3.
@@ 3.1 Segmentation Priors from Frozen LMM
Vision-Language Sequence. A typical build of a LMM  [^2] In this paper, the term ‘multimodal’ stands for vision and language modalities. comprises an image encoder $f_{v}$ (e.g., CLIP [58]  [^3] The image encoder might be any vision model that is pre-trained on image-text pairs. We use the classic term ‘CLIP’ in this paper to represent all such models for brevity.), a vision-language projector $f_{p}$, and a LLM, denoted as $f_{\mathrm{llm}}$. The inputs to an LMM are usually an image ${\mathbf{X}}_{v}\in{\mathbb{R}}^{3\times H\times W}$ and the associated text ${\mathbf{X}}_{t}$. The input image is first encoded by the vision encoder $f_{v}$ and then mapped to the input space of the LLM by the projector $f_{p}$:
${\mathbf{Z}}_{v}=f_{p}(\texttt{Flatten}(f_{v}({\mathbf{X}}_{v})))\in{\mathbb{R}}^{hw\times d},$
@@
where $h$ and $w$ are the height and width of projected feature maps via $f_{v}$. The Flatten operation unfolds the 2-D image feature map to a 1-D sequence. The constant $d$ is the hidden state dimension of the LLM. Likewise, the text input is first encoded as discrete tokens and then mapped to text embeddings:
${\mathbf{Z}}_{t}=\texttt{Embed}(\texttt{Tokenize}({\mathbf{X}}_{t}))\in{\mathbb{R}}^{L\times d},$
@@
where $L$ denotes the length of text embeddings. The visual-language sequence input to the LLM is a concatenation of image and text embeddings: ${\mathbf{Z}}=\{{\mathbf{Z}}_{v},{\mathbf{Z}}_{t}\}\in{\mathbb{R}}^{(hw+L)\times d}$.
@@
Table 1: The main evaluation results on question-answering benchmarks, referring expression segmentation (RES) benchmark and panoptic narrative grounding (PNG) benchmark. LLaVAW: LLaVA-In-the-Wild. LLaVA-1.6 and MGM-HD take high-resolution image inputs. LLaVA-1.6-M-7B means the model is based on Mistral-7B [22]. GLaMM-FS-7B means we use the ‘FullScope’ version of GLaMM.
@@
MME MMBench MMVet LLaVAW RefCOCO RefCOCO+ RefCOCOg All Thing Stuff